OpenStudy (anonymous):

solve the equation 8x^3+12x^2+6X+1=0

OpenStudy (anonymous):

factor your polynomial, note that\[8x^3+12x^2+6x+1=8x^3+1+12x^2+6x\]\[=(2x)^3+1+6x(2x+1)\]you may make use of the fact that\[a^3+b^3=(a+b)(a^2-ab+b^2)\]for factoring \((2x)^3+1\)